OpenStudy (anonymous):

Which of the following is a solution to the equation log 4 x + log 4 (x – 3) = 1 ? x = 6/5 x = 2 x = –3 x = 4

OpenStudy (anonymous):

Use this on left hand side: \[\large \log(a) + \log(b) = \log(ab)\]

OpenStudy (anonymous):

is 4 the base ??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[\log_4(x) + \log_4(x-3) = 1 \implies \log_4(x(x-3)) = \log_4(4)\] \[x(x-3) = 4\] Can you proceed further.???

OpenStudy (anonymous):

See: \[x(x-3) = 4 \implies 4(4 - 3) = 4 \times 1\] So x = 4 is the solution...

OpenStudy (anonymous):

so its four?

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

kk :) thanksssss

OpenStudy (anonymous):

Welcome dear..

OpenStudy (jiteshmeghwal9):

for learning the properties of Logarithms look at this http://openstudy.com/users/jiteshmeghwal9#/updates/5002ba39e4b0848ddd66b602